# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ 3.5 Confidence Intervals for LOORA-HT
We now describe how to estimate the variance of the LOORA-HT estimator and construct corresponding confidence intervals. We treat $\widehat{\tau}_{\textup{LHT}}$ as a two-step estimator and show that the consistency of the second step does not depend on the consistency of the first-step parameter estimates. This implies that our variance estimator remains valid even when the first-step ridge coefficients $\widehat{\boldsymbol{\beta}}_{\lambda}$ are not estimated consistently. To estimate the variance in the second step, we employ the Huber–White (or “sandwich”) estimator of asymptotic variance.
Recall that LOORA-HT estimates the average treatment effect as
$\widehat{\tau}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}z_{i}\!\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}}{q_{i}}\right)=\frac{1}{n}\sum_{i=1}^{n}z_{i}\!\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\right),$
where the second equality follows from Lemma˜B.1. Hence $\widehat{\tau}_{\textup{LHT}}$ can be interpreted as the coefficient from regressing the vector
$\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{q_{i}(1-\widetilde{h}_{\lambda ii})}\right)_{i\in[n]}\quad\text{on}\quad{\bm{z}}=(z_{1},\ldots,z_{n})^{\top}.$
This representation makes it clear that $\widehat{\tau}_{\textup{LHT}}$ is a two-step estimator: in the first step, we obtain $\widehat{\boldsymbol{\beta}}_{\lambda}$ from a ridge regression of outcomes on covariates, and in the second step, we estimate the coefficient from a regression of the residualized outcomes on the treatment indicator. Although in practice we compute only the leave-one-out coefficients $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$, the residuals computed using $\widehat{\boldsymbol{\beta}}_{\lambda}$ and scaling with $(1-\widetilde{h}_{\lambda ii})^{-1}$ are equivalent by Lemma˜B.1.
Importantly, the consistency of $\widehat{\tau}_{\textup{LHT}}$ does not depend on whether $\widehat{\boldsymbol{\beta}}_{\lambda}$ consistently estimates the population limit $\boldsymbol{\beta}^{*}_{\lambda}$. Indeed, by construction,
${\mathbb{E}}\!\left[z_{i}\!\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}}{q_{i}}\right)\right]=y_{i}^{(1)}-y_{i}^{(0)},$
so that $\widehat{\tau}_{\textup{LHT}}$ remains unbiased for $\tau$ and hence consistent under random assignment.
Huber–White variance and confidence intervals.
For this univariate regression, the residual for unit $i$ is
$\widehat{r}_{i}:=\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{q_{i}(1-\widetilde{h}_{\lambda ii})}-z_{i}\,\widehat{\tau}_{\textup{LHT}}.$
Let ${\bm{Z}}=(z_{1},\ldots,z_{n})^{\top}$ and $\widehat{{\bm{R}}}=(\widehat{r}_{1},\ldots,\widehat{r}_{n})^{\top}$. The Huber–White (HC0) variance estimator for the slope is
$\widehat{V}_{\textup{HW}}=({\bm{Z}}^{\top}{\bm{Z}})^{-1}{\bm{Z}}^{\top}\operatorname{diag}(\widehat{{\bm{R}}}\odot\widehat{{\bm{R}}}){\bm{Z}}({\bm{Z}}^{\top}{\bm{Z}})^{-1}=\frac{\sum_{i=1}^{n}z_{i}^{2}\,\widehat{R}_{i}^{2}}{(\sum_{i=1}^{n}z_{i}^{2})^{2}}.$
Since $z_{i}\in\{\pm 1\}$, this simplifies to $\widehat{V}_{\textup{HW}}=n^{-2}\sum_{i=1}^{n}\widehat{r}_{i}^{2}$. This estimator consistently estimates the asymptotic variance derived in Theorem 4 under mild moment conditions, and the “sandwich” form adjusts for the heteroskedasticity induced by unit-specific $p_{i}$ and the first-step estimation.
Finally, a $(1-\alpha)$ confidence interval for the average treatment effect is
$\widehat{\tau}_{\textup{LHT}}\;\pm\;z_{1-\alpha/2}\sqrt{\widehat{V}_{\textup{HW}}},$
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$ quantile of the standard normal distribution. This interval is asymptotically valid under the regularity conditions stated in Section 3.3.
@@ 4 Regression Adjustment for Difference-in-Means Estimator
We now consider the difference-in-means (DM) estimator under complete random assignment, where exactly $n_{T}$ units are assigned to treatment and $n_{C}=n-n_{T}$ to control. Throughout this subsection, the signal vector used in the Horvitz–Thompson analysis is redefined. Specifically,
$\widetilde{{\boldsymbol{\mu}}}\;:=\;n_{C}\,{\bm{y}}^{(1)}\;+\;n_{T}\,{\bm{y}}^{(0)},$
and all formulas below are written in terms of $\widetilde{{\boldsymbol{\mu}}}$ rather than the $\boldsymbol{\mu}$} used earlier.
@@
where $S_{n}^{2}(\cdot)$ denotes finite-population variances computed with the (n1)(n-1) divisor (see, e.g., imbens2015causal). Equivalently—and most convenient for our regression-adjustment analysis—this variance admits the representation
${\rm Var}\!\big(\widehat{\tau}_{\textup{DM}}\big)\;=\;\frac{1}{\,n_{T}\,n_{C}\,n\,(n-1)\,}\;\big\|\,\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}}\,\big\|_{2}^{2},$ (12)
where $\overline{\widetilde{{\boldsymbol{\mu}}}}:=(n^{-1}\sum_{i=1}^{n}\widetilde{\mu}_{i})\boldsymbol{1}$.
For covariate-adjusted outcomes $\breve{y}_{i}({\bm{b}})=y_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}$, with fixed ${\bm{b}}\in\mathbb{R}^{k}$, the adjusted DM estimator
$\frac{1}{n_{T}}\sum_{i=1}^{n}d_{i}\,\breve{y}_{i}({\bm{b}})\;-\;\frac{1}{n_{C}}\sum_{i=1}^{n}(1-d_{i})\,\breve{y}_{i}({\bm{b}})$
remains unbiased for $\tau$, and its variance is
${\rm Var}\!\big(\widehat{\tau}_{\textup{DM}}({\bm{b}})\big)\;=\;\frac{1}{\,n_{T}\,n_{C}\,n\,(n-1)\,}\;\big\|\,\big(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}}\big)-\overline{\big(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}}\big)}\,\big\|_{2}^{2},$ (13)
where $\overline{(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}})}:=(n^{-1}\sum_{i=1}^{n}(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}))\boldsymbol{1}$. Therefore, the minimum attainable variance is
$\frac{1}{\,n_{T}\,n_{C}\,n\,(n-1)\,}\;\min_{{\bm{b}}\in\mathbb{R}^{k}}\big\|\,({\bm{X}}-\overline{{\bm{X}}})\,{\bm{b}}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\,\big\|_{2}^{2}.$ (14)
However, since $\widetilde{{\boldsymbol{\mu}}}$ is not observable—each unit reveals only one potential outcome—this variance cannot be achieved in practice. In what follows, we construct an estimable surrogate with the correct conditional expectation and develop a leave-one-out regression-adjusted DM estimator (LOORA-DM), together with its finite-sample variance expression and asymptotic properties under complete random assignment.
@@ 4.1 Leave-One-Out Regression-Adjusted Difference-in-Means
Similar to LOORA-HT, our proposed leave-one-out regression-adjusted difference-in-means (LOORA-DM) estimator (Algorithm˜2) computes a distinct coefficient vector $\bm{b}$} for each unit $i$. A key difference, however, is that the regressed vector $\widetilde{{\bm{y}}}$ also differs across units. Under complete random assignment, the treatment assignments are not independent, so the regressed vector must be scaled differently depending on whether unit $i$ belongs to the treatment or control group in order to ensure that it has the correct expectation.
@@
3:Calculate ${\bm{f}}_{T}\in{\mathbb{R}}^{n}$ with $f_{Ti}=\frac{1}{n_{T}(n_{T}-1)}$ if $d_{i}=1$, and $f_{Ti}=\frac{1}{n^{2}_{C}}$ otherwise.
4:Calculate ${\bm{f}}_{C}\in{\mathbb{R}}^{n}$ with $f_{Ci}=\frac{1}{n_{T}^{2}}$ if $d_{i}=1$, and $f_{Ci}=\frac{1}{n_{C}(n_{C}-1)}$ otherwise.
@@
$S=S+v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}).$
@@
$\widehat{\tau}=S.$
More precisely, for each unit $i$, we define
@@
$f^{(-i)}_{j}=\begin{cases}\dfrac{1}{n_{T}(n_{T}-1)},&\text{if }d_{j}=1,\\[10.00002pt] \dfrac{1}{n_{C}^{2}},&\text{if }d_{j}=0,\end{cases}\qquad\text{if }d_{i}=1,$
and
$f^{(-i)}_{j}=\begin{cases}\dfrac{1}{n_{T}^{2}},&\text{if }d_{j}=1,\\[10.00002pt] \dfrac{1}{n_{C}(n_{C}-1)},&\text{if }d_{j}=0,\end{cases}\qquad\text{if }d_{i}=0.$
The coefficient vector $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ is then obtained by regressing $\widetilde{{\bm{y}}}^{(-i)}_{-i}$ on ${\bm{X}}_{-i}$ using ridge regression. Finally, the outcome of unit $i$ is adjusted by ${\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$.
In the next subsection, we establish that LOORA-DM is unbiased for the average treatment effect and derive an exact expression for its variance.
4.2 The Variance of LOORA-DM
We now characterize the finite-sample variance of the leave-one-out regression-adjusted difference-in-means (LOORA-DM) estimator. The terms in the variance are more intricate here because, under complete random assignment, treatment assignments are not independent. Unbiasedness is also slightly more delicate to verify than in the Horvitz–Thompson setting, but it ultimately follows from the fact that $\widetilde{{\bm{y}}}^{(-i)}$ is constructed so that ${\mathbb{E}}\!\left[\widetilde{{\bm{y}}}^{(-i)}_{-i}\right]=\frac{\widetilde{{\boldsymbol{\mu}}}_{-i}}{n}$ for each $i$.
@@ 4.2 The Variance of LOORA-DM
Under complete random assignment, the estimator obtained from Algorithm˜2 is unbiased and its variance is given by
$\frac{1}{n(n-1)n_{T}n_{C}}\sum_{i=1}^{n}\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda ii}}-\frac{1}{n}\sum_{j=1}^{n}\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda jj}}\right)^{2}$
$\quad-\frac{2}{n^{2}(n-1)n_{T}n_{C}}\left(\sum_{\begin{subarray}{c}i,j\in[n]:\\ i\neq j\end{subarray}}\frac{(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})\,h_{\lambda ij}\,\widetilde{\mu}_{i}}{(1-h_{\lambda ii})(1-h_{\lambda jj})}-\frac{1}{n-2}\sum_{\begin{subarray}{c}k\in[n]:\\ k\neq i,j\end{subarray}}\frac{(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})\,h_{\lambda jk}\,\widetilde{\mu}_{k}}{(1-h_{\lambda ii})(1-h_{\lambda jj})}\right)$
@@
$\widetilde{{\boldsymbol{\mu}}}$ $=n_{C}{\bm{y}}^{(1)}+n_{T}{\bm{y}}^{(0)},\qquad\boldsymbol{\beta}_{\lambda}=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\widetilde{{\boldsymbol{\mu}}}$,
@@
Entries of $\boldsymbol{Q}$ depend only on $n_{C}$, $n_{T}$, and the entries of the hat matrix $\boldsymbol{H}_{\lambda}$; its explicit form is provided in the appendix.
The first line in the variance expression can be viewed as the squared Euclidean norm of a vector after subtracting its sample mean. Specifically, it is the norm of
$\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda ii}}\right)_{i\in[n]}\;-\;\frac{1}{n}\sum_{j=1}^{n}\left(\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda jj}}\right)\mathbf{1},$
scaled by the factor $n(n-1)n_{T}n_{C}]^{-1}$. The denominators $1-h_{\lambda ii}$ provide the leverage correction arising from leave-one-out adjustment. By contrast, in the Horvitz–Thompson case one must recenter the matrix and add an intercept explicitly to produce the analogous centering.
The second line collects cross-unit terms governed by the off-diagonal leverages $h_{\lambda ij}$. These cross-unit contributions did not appear in the Horvitz–Thompson analysis because units are independent under simple random assignment, whereas here the fixed treatment counts induce dependence across units under complete random assignment. The final quadratic form in $\boldsymbol{Q}$ bundles additional corrections that depend only on $(n_{T},n_{C})$ and the leverage structure encoded in $\boldsymbol{H}_{\lambda}$. For a fixed number of covariates, the first term scales as $O(1/n)$, while both the cross-unit term and the $\boldsymbol{Q}$-term scale as $O(1/n^{2})$; in particular, the second and third terms vanish asymptotically.
Taken together, Theorem 5 shows that the variance of LOORA-DM is dominated by the centered dispersion of the leverage-adjusted signal and that dependence induced by complete random assignment contributes only second-order terms. This decomposition will be useful when establishing asymptotic distributional results and efficiency comparisons later in the section.

### v2

@@ 3.5 Confidence Intervals for LOORA-HT
The next theorem presents a variance estimator for $\widehat{\tau}_{\textup{LHT}}$.
Theorem 4.
Suppose Assumptions˜1, 2 and 3 hold. Let $\lambda\geq 0$ be fixed,
$\widehat{s}_{i}$ $=z_{i}\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}}{q_{i}}\right)-\,\widehat{\tau}_{\textup{LHT}},~~\ \widehat{V}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\widehat{s}_{i}^{2},~~\text{and}$
$V_{n}$ $=\frac{1}{n}\left(\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}^{2}+\sum_{i=1}^{n}(\tau_{i}-\tau)^{2}\right),$
where $\tau_{i}=y^{(1)}_{i}-y^{(0)}_{i}$. If $V_{n}$ is bounded away from zero, then
$\frac{\widehat{V}_{\textup{LHT}}}{V_{n}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}1.$
As is generally the case in all finite-population inference (see, e.g., Abadie et al., 2020), $\widehat{V}_{\textup{LHT}}$ is consistent under homogeneous individual treatment effects and conservative when treatment effects are heterogeneous. We construct confidence intervals using the estimate of standard deviation as $\sqrt{\widehat{V}_{\textup{LHT}}/n}$ and the appropriate quantiles of the standard normal distribution.
@@ 4 Regression Adjustment for Difference-in-Means Estimator
This section considers the difference-in-means (DM) estimator under complete random assignment, where exactly $n_{T}$ units are assigned to treatment and $n_{C}=n-n_{T}$ to control.
@@
where $S_{n}^{2}(\cdot)$ denotes finite-population variances computed with an n1n-1 divisor (see, e.g., Imbens and Rubin (2015, Chapter 6, Appendix A)). Equivalently—and most convenient for our regression-adjustment analysis—this variance admits the representation
${\rm Var}\!\big(\widehat{\tau}_{\textup{DM}}\big)\;=\;\frac{1}{\,n\,(n-1)\,}\;\big\|\,\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}}\,\big\|_{2}^{2},$
where
$\widetilde{{\boldsymbol{\mu}}}\;=\;\sqrt{\frac{n_{C}}{n_{T}}}\,{\bm{y}}^{(1)}\;+\;\sqrt{\frac{n_{T}}{n_{C}}}\,{\bm{y}}^{(0)}.$ (8)
For a fixed vector ${\bm{b}}\in\mathbb{R}^{k}$, the covariate-adjusted DM estimator
$\widehat{\tau}_{\textup{DM}}({\bm{b}})=\frac{1}{n_{T}}\sum_{i=1}^{n}d_{i}\,(y_{i}-\eta^{-1}{\bm{x}}_{i}^{\top}{\bm{b}})\;-\;\frac{1}{n_{C}}\sum_{i=1}^{n}(1-d_{i})\,(y_{i}-\eta^{-1}{\bm{x}}_{i}^{\top}{\bm{b}}),$
where $\eta=\sqrt{n_{C}/n_{T}}+\sqrt{n_{T}/n_{C}}$, is unbiased for $\tau$, and its variance is equal to
${\rm Var}\!\big(\widehat{\tau}_{\textup{DM}}({\bm{b}})\big)\;=\;\frac{1}{\,n\,(n-1)\,}\;\big\|\,\big(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}}\big)-\overline{\big(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}}\big)}\,\big\|_{2}^{2},$ (9)
with $\overline{(\widetilde{{\boldsymbol{\mu}}}-{\bm{X}}{\bm{b}})}=(n^{-1}\sum_{i=1}^{n}(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}))\boldsymbol{1}$ . Therefore, the minimum variance of $\widehat{\tau}_{\textup{DM}}({\bm{b}})$ is equal to
$\frac{1}{\,n\,(n-1)\,}\;\min_{{\bm{b}}\in\mathbb{R}^{k}}\left\|\,({\bm{X}}-\overline{{\bm{X}}})\,{\bm{b}}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\,\right\|_{2}^{2}.$ (10)
The vector $\widetilde{{\boldsymbol{\mu}}}$ is not observed, so the variance in (10) cannot be attained in practice. The next section introduces an estimable surrogate and uses it to define a leave-one-out regression-adjusted DM estimator (LOORA-DM).
@@ 4.1 Leave-One-Out Regression-Adjusted Difference-in-Means
Similar to LOORA-HT, our proposed leave-one-out regression-adjusted difference-in-means (LOORA-DM) estimator (Algorithm˜2) computes a distinct coefficient vector $\bm{b}$} for each unit $i$. A key difference, however, is the dependent variable used in the regression to estimate the adjustment coefficients. This regression must be modified to account for the dependence in treatment assignments induced by complete randomization.
@@
3:Calculate ${\bm{f}}_{T}\in{\mathbb{R}}^{n}$ with $f_{Tj}=\frac{1}{n_{T}(n_{T}-1)}$ if $d_{j}=1$, and $f_{Tj}=\frac{1}{n^{2}_{C}}$ otherwise.
4:Calculate ${\bm{f}}_{C}\in{\mathbb{R}}^{n}$ with $f_{Cj}=\frac{1}{n_{T}^{2}}$ if $d_{j}=1$, and $f_{Cj}=\frac{1}{n_{C}(n_{C}-1)}$ otherwise.
@@
$S=S+v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda})$.
@@
$\widehat{\tau}_{\textup{LDM}}=S.$
For each unit $i$, we define
@@
$f^{(-i)}_{j}=\begin{cases}\dfrac{1}{n_{T}(n_{T}-1)}&\text{if }d_{j}=1,\\[10.00002pt] \dfrac{1}{n_{C}^{2}}&\text{if }d_{j}=0,\end{cases}$
if $d_{i}=1$, and
$f^{(-i)}_{j}=\begin{cases}\dfrac{1}{n_{T}^{2}}&\text{if }d_{j}=1,\\[10.00002pt] \dfrac{1}{n_{C}(n_{C}-1)}&\text{if }d_{j}=0,\end{cases}$
if $d_{i}=0$.
The coefficient vector $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ is then obtained by regressing $\widetilde{{\bm{y}}}^{(-i)}_{-i}$ on ${\bm{X}}_{-i}$ using ridge regression. Finally, the outcome of unit $i$ is adjusted by ${\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$,
$\widehat{\tau}_{\textup{LDM}}=\frac{1}{n_{T}}\sum_{i=1}^{n}d_{i}\,(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda})\;-\;\frac{1}{n_{C}}\sum_{i=1}^{n}(1-d_{i})\,(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda})$.
4.2 The Bias and Variance of LOORA-DM
This section establishes that LOORA-DM is exactly unbiased for the average treatment effect and derives an exact expression for its finite-sample variance. The expression of the variance is more intricate for LOORA-DM than for LOORA-HT because complete randomization induces dependence between treatment assignments. Unbiasedness is also slightly more delicate to verify than in the LOORA-HT case, but it ultimately follows from the fact that $\widetilde{{\bm{y}}}^{(-i)}$ is constructed so that ${\mathbb{E}}[\widetilde{{\bm{y}}}^{(-i)}_{-i}]=(\sqrt{n_{c}n_{T}}/n)\widetilde{{\boldsymbol{\mu}}}_{-i}$ for each $i$.
@@ 4.2 The Bias and Variance of LOORA-DM
Under complete random assignment, LOORA-DM is unbiased with variance equal to
$\frac{1}{n(n-1)}\sum_{i=1}^{n}\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda ii}}-\frac{1}{n}\sum_{j=1}^{n}\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{1-h_{\lambda jj}}\right)^{2}$
$\quad-\frac{2}{n^{2}(n-1)}\sum_{\begin{subarray}{c}i,j\in[n]:\\ i\neq j\end{subarray}}\left(\frac{(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})\,h_{\lambda ij}\,\widetilde{\mu}_{i}}{(1-h_{\lambda ii})(1-h_{\lambda jj})}-\frac{1}{n-2}\sum_{\begin{subarray}{c}k\in[n]:\\ k\neq i,j\end{subarray}}\frac{(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})\,h_{\lambda jk}\,\widetilde{\mu}_{k}}{(1-h_{\lambda ii})(1-h_{\lambda jj})}\right)$
@@
$\boldsymbol{\beta}_{\lambda}$ $=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\widetilde{{\boldsymbol{\mu}}}$,
@@
The entries of $\boldsymbol{Q}$ depend only on $n_{C}$, $n_{T}$, and the entries of the hat matrix $\boldsymbol{H}_{\lambda}$; its explicit form is provided in the appendix.
The first line in the variance expression is the finite-population variance of $(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})/(1-h_{\lambda ii})$ rescaled by $1/n$. The $1-h_{\lambda ii}$ in the denominator provides a leverage correction that generates from leave-one-out adjustment. The second line of the variance expression in Theorem˜5 collects cross-unit terms that include the off-diagonal leverages $h_{\lambda ij}$. These cross-unit contributions did not appear in the LOORA-HT analysis because simple random assignment makes assignments independent across units, whereas complete random assignment induces dependence across units. The matrix $\boldsymbol{Q}$ bundles additional corrections that depend only on $(n_{T},n_{C})$ and the leverage structure encoded in $\boldsymbol{H}_{\lambda}$. For a fixed number of covariates, the first term scales as $O(1/n)$, while both the cross-unit term and the $\boldsymbol{Q}$-term scale as $O(1/n^{2})$. [^2] To see why the cross-unit term in the result of Theorem 5 is $O(1/n^{2})$, apply Cauchy-Schwarz inequality and note that $(\sum_{\begin{subarray}{c}i,j\in[n]:i\neq j\end{subarray}}h_{\lambda ij}^{2})^{0.5}\leq\|\boldsymbol{H}\|_{F}\leq\sqrt{k}$.
Theorem 5 shows that the variance of LOORA-DM is dominated by the centered dispersion of the leverage-adjusted signal and that dependence induced by complete random assignment contributes only second-order terms. This decomposition is useful to establish asymptotic distributional results and efficiency comparisons.